OpenStudy (anonymous):

An initial population of 745 quail increases at an annual rate of 16%. Write an exponential function to model the quail population. What will the approximate population be after 4 years?

OpenStudy (anonymous):

@agent0smith can you help plzzzz?:( i have to hurry and get this done D:

OpenStudy (anonymous):

A = P (1 + r)^t solve

OpenStudy (anonymous):

hmmm im not really sure, but ill try A=P1^t+Pr^t ?? @sourwing

OpenStudy (anonymous):

@jdoe0001 please hellpppp

OpenStudy (nikato):

Ok, so the formula for exponential is y=a(1+r)^x

OpenStudy (nikato):

Where a=initial amount R=rate And x=time

OpenStudy (agent0smith):

\[\Large A = P (1 + r)^t\] so r = 0.16, P = 745\[\LARGE A = 745 (1+0.16)^t\] Do things in brackets FIRST. Not as you did above.

OpenStudy (nikato):

And t=4^

OpenStudy (anonymous):

so would it be A=745(1.16)^t

OpenStudy (agent0smith):

A = P (1 + r)^t A=P1^t+Pr^t DO NOT DO THIS. You cannot distribute when there's an exponent, that's not following order of operations.

OpenStudy (nikato):

Yes, that's correct. And t=4 So ur next step would be 1.26^4

OpenStudy (nikato):

1.16*

OpenStudy (nikato):

According to PEMdAS

OpenStudy (anonymous):

so would it be A=1348.92??

OpenStudy (anonymous):

as the final answer?

OpenStudy (agent0smith):

Yes

OpenStudy (anonymous):

ok so that would be the approximate population after 4 years? so the exponential function would be A=745(1.16)^t ???

OpenStudy (nikato):

Yup

OpenStudy (anonymous):

what is the exact population? because you can't have a decimal for people lol

OpenStudy (nikato):

I guess u can round up to 1349

OpenStudy (agent0smith):

Round up or down, it's an approximation so either is fine. Round up is usually better for real things.

OpenStudy (agent0smith):

or just say ~1350

OpenStudy (agent0smith):

"what is the exact population?" note the question doesn't ask for exact population.

OpenStudy (anonymous):

ok! and is the exponential function that i wrote up above correct?

OpenStudy (nikato):

Yes, it is correct

OpenStudy (anonymous):

are you positive?(:

OpenStudy (agent0smith):

lol I gave you a medal in the other question because it was correct!

OpenStudy (anonymous):

ok thanks guyssss! (: